OpenStudy (anonymous):

2x-4y=6 x-2y=3 substitution

OpenStudy (anonymous):

are we trying to solve for both x and y? and do we have any given values?

OpenStudy (anonymous):

2x-4y=6 lets take 2 common from both the sides => 2 ( x -2y) = 2(3) => x - 2y = 3 and the other equation that we have is x-2y=3 since the equations are equal It implies they lie on the same line and hence infinite number of solutions exist